OpenStudy (anonymous):

Question about partial differentiation: f(x,y)=(e^x)(siny) find values of x and y when fx=0 and fy=0 simultaneously

hartnn (hartnn):

what u got as fx=0 ?

OpenStudy (anonymous):

(e^x)(siny)

hartnn (hartnn):

e^x can't be 0, so sin y=0

OpenStudy (anonymous):

fy=(e^x)(cosy)

hartnn (hartnn):

similarly, for fy=0, u get cos y=0

OpenStudy (anonymous):

what about if i divide fx and fy so that e^x is eliminated and left tany=0?

hartnn (hartnn):

tan y =0 will imply that sin y=0 and cos y NOT =0 which is not true.

hartnn (hartnn):

thats because you cannot divide by 0(e^x cos x=0)

hartnn (hartnn):

got that ?

hartnn (hartnn):

instead of dividing, why don't u multiply ? and use sin 2x = 2 sin x cos x

OpenStudy (anonymous):

oh,ok.I m trying nw,thx ^^